OpenStudy (anonymous):

The ion product of water: varies with pH is only applicable to neutral water remains constant regardless of the pH ranges between 1.0 and 10-14 Help please this the last question and it determines if I pass or not D: Please Help..

OpenStudy (anonymous):

@timo86m help if you can please

OpenStudy (frostbite):

The equation for self-ionization of water: \[\LARGE K_{W}=\left[ H ^{+} \right] ~ \left[ OH ^{-} \right]=1.0*10^{-14}\] Kw is a constant, but temperature dependent and last equal sign is true at 25 °C. H+ and OH- are variables. What state is true based upon that equation?